Could someone please show me how to find the intercepts and then graph the equation?
1.9x-1.1y=4.18

Answer by mananth(16946)

1.9x-1.1y=4.18
1.1 y = 1.9x-4.18
multiply by 10
11y= 19x-41.8
/11
y=19x/11 -41.8/11
y=19x/11 - 3.8
when x = 0 y= -3.8 (0,-3.8)
when y = 0
19x/11= 3.8
19x=3.8*11
19x=41.8
x= 41.8/19
x=2.2
(2.2,0)
You have the two intercept points (0,-3.8)& (2.2,0)
Plot these two points and draw a line passing through these two points.